Long-term history has shown that 65% of all elected offices in a rural county have been won by Republican candidates. This year there are 5 offices up for public election in the county Let r be the number of public offices won by Republicans.
a) Find P(r) for r=0,1,2,3,4, and 5
b) Make a histogram for the r probability distribution.
c) What is the expected number of Republicans who will win office in the coming election?
d) What is the standard deviation of r?

Answer by stanbon(75887)

Long-term history has shown that 65% of all elected offices in a rural county have been won by Republican candidates. This year there are 5 offices up for public election in the county Let r be the number of public offices won by Republicans.
a) Find P(r) for r=0,1,2,3,4, and 5
General Formula: P(x = r) = 5Cr*0.65^r*0.35^(5-r)
Eg: P(x=3) = 5C3*0.65^3*0.35^2 = 10*0.2746*0.1225 = 0.3364...
-----------------------------------------------------------
b) Make a histogram for the r probability distribution.
I'll let you do that.
What is the expected number of Republicans who will win office in the coming election?
E(x) = 0*P(0)+1*P(1)+2*P(2)+3*P(3)+...
----------------------------------------
d) What is the standard deviation of r?
It's binomial so it's sqrt(npq) = sqrt(5*0.65*0.35) = 1.0665...
Cheers,
Stan H.
Cheers,
Stan H.